2 identical sphere having charges +10 micro coulomb and -90 micro coulomb attract each other with a force F. If they are kept in contact and then separated by the same distance, the new force between them will be..
a)F/6 b)16F c) 16F/9 d)9F    (Please tell me the steps!)


Thank u! 

Grade:12

3 Answers

shiddhant bhattacharya
11 years ago

is the ans c?

Rohan Nambiar
11 years ago

yeh!but how did u get it?can u show me the method?

Subhasmita pradhan
7 years ago
When they are touched charge will be equally distributed.so(10+(-90))/2=-40So new force will be F`=K(-40)^2/r^2Initially the force wasF=10×-90/r^2Deviding both the forcesF`=16F/9